OpenStudy (anonymous):

using a scale factor of 1 inch =5 feet, what are the blue print diminsions of the bed room that is 12 feet x 18 feet ? a)2.2 inch x 3.3 inch b.)2.4 inch x 3.6 inch c.)2.8 inch x 4.2inch d.)2.9inch x 5.0 inch

OpenStudy (anonymous):

i would say a

OpenStudy (texaschic101):

1/5 = x/12 -- 1 in to 5 ft = x in to 12 ft cross multiply 5x = 12 x = 12/5 x = 2 2/5 in. wich equals 2.4 in 1/5 = x/18 -- 1 in to 5 ft = x in to 18 ft cross multiply 5x = 18 x = 18/5 in. which equals 3.6 in answer is : B